OpenStudy (anonymous):

Combine the like terms to simplify the expression: 5 + 7 + 15xy2 + 9y3 + 10x –13xy2 – 3x – 8y3 A. y3 + 2xy2 + 10x + 1 B. y6 + 2x2y2 + 10x + 12 C. y3 + 2xy2 + 7x + 12 D. y3 –2xy2 + 10x + 15

OpenStudy (lynfran):

which option do u think it is?

OpenStudy (anonymous):

D?

OpenStudy (lynfran):

no

OpenStudy (lynfran):

5+7=12

OpenStudy (lynfran):

15xy2-13xy2=2xy2

OpenStudy (lynfran):

9y3-8y3=y3

OpenStudy (lynfran):

10x-3x=7x

OpenStudy (lynfran):

now which option is correct

OpenStudy (anonymous):

c?

OpenStudy (lynfran):

yes correct

OpenStudy (anonymous):

thank you so much :D

OpenStudy (lynfran):

welcome